The left and right page numbers of an open book are two consecutive integers whose sum is 209. Find these page numbers.


1
Expert's answer
2021-01-29T16:29:31-0500

Left page = x


Right page = x + 1


x + (x + 1) = 209


2 x + 1 = 209


2 x = 208


x = "\frac{208}{2}" =104


Left page = 104


Right page = 105